OpenStudy (erniellyd28):

at 10 a.m a plane leaves boston, massachusetts, for seattle, washington,a distance of 3000 mi. one hour later a plane leaves seattle for boston. Both planes are traveling at a speed of 500 mph. how many hours after after the plane leaves seattle will the planes pass each other

pooja195 (pooja195):

. Let t = time of travel of plane leaving Boston then t -1 = time of travel of plane leaving Seatle Now the sum of thier distances = 3000 so we have 500t + 500(t-1) = 3000 then 500t + 500t - 500 = 3000 then 1000t = 3500 and t = 3.5 hrs. Then t-1 = 2.5 hours.

OpenStudy (erniellyd28):

do i have to do any thing any thing else as afar as solving after 2.5 or thats the answer

pooja195 (pooja195):

yep answer

OpenStudy (erniellyd28):

thank u so very much for all ur help